A river descends 15 m through rapids. The speed of the water is 3.2 m/s upon entering the rapids and 13 m/s upon leaving. What percentage of the gravitational P.E of the water- Earth system is transferred to the kinetic energy during the descent ? 

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear Sanchit

Max potential energy that can be converted in to kinetic energy is =m * g*15=150*m


but the tranfered potential energy in to kinetic energy is =1/2 *m (132-3.22)=79.38 *m


so percentage change is ={79.38 *m/150*m} *100

                                  =52.92 %


Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin